OpenStudy (anonymous):

verify the identity 1/sin(ø-1) + 1/sin(ø+1) = -2tanøsecø

OpenStudy (calculusfunctions):

Work with the left side. Step 1: Add the fractions

OpenStudy (anonymous):

im confused how to even start it. like will the left side be 2sinø by adding them?

OpenStudy (anonymous):

i mean 2/sinø

OpenStudy (calculusfunctions):

How do you add fractions @biochems2k ? You first find the least common denominator, correct?

OpenStudy (anonymous):

correct, so multiply both of them by sinø - 1

OpenStudy (calculusfunctions):

\[Prove: \frac{ 1 }{ \sin Φ-1 }+\frac{ 1 }{ \sin Φ+1 }=2\tan Φ\sec Φ\]\[L.S.=\frac{ 1 }{ \sin Φ-1 }+\frac{ 1 }{ \sin Φ+1 }\]Least common denominator is (sin Φ - 1)(sin Φ + 1). Understand @biochems2k ?

OpenStudy (calculusfunctions):

Thus what is the sum of the two fractions?

OpenStudy (calculusfunctions):

If you don't know and would like me to show you the first step, please do not be embarrassed to say so.

OpenStudy (anonymous):

sorry, i was doing the work as you explained. after doing all that i at the end 2/sin^2ø - 1. am i correct?

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

@calculusfunctions im actually stuck now. my answer didnt look correct

OpenStudy (anonymous):

@calculusfunctions after doing my math again i got 2sinø/sinø-1